User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!

Meeting 25th May

Chair: Jesse Cleaning: Wut

Present: Thanks!

Apologies: None

Absent: None

General Issues:

* Hb12 Rfpic - Hardware arrived, pending fix. 
* Hb12 Drives - Only master drive was running, HMI interface showed that One azimuth drive was down, it still reads as tracking in operin. Reset resolved this, but there were no alarms, so KEEP AN EYE OUT.  
   * Drives wouldn't go into operating, a restart was necessary (Best to do a restart before experiment starts).  
* Ke12 - Pointing check took a long time as Ke would continue to lose onsource.
* Fringes -- Seem to work infrequently, this is not likely to be a data quality issue but please check power levels just to be safe. 
* if power levels take a long time to get to target (5--10min) due to interference in s-band. x-band is all tickity-boo though. 
* Yg being a little too aggressive wind windstows (7--10 km/h). Reset the anemometer (y) 
   * Can check the wind-speed reading on the cameras.  
* Ke mk5 connection error (connecting fine, except that occasionally response is slow enough to trigger an alarm)  
* Jonny? m8? Cmon.  

Handover Issues:

/home/www/auscope/opswiki/data/pages/operations/agenda2017_05_11.txt · Last modified: 2017/05/25 03:39 by Jesse Swan